OpenStudy (anonymous):

write a recursive formula for the sequence 7, 4, 1, -2, -5, ...then find the next term.

OpenStudy (anonymous):

@Hero

hero (hero):

Hang on for a minute while I solve this

hero (hero):

Okay, I have it

hero (hero):

So the first thing you should do is create a chart of values (n , r) (1, 7), (2, 4), (3, 1), (4, -2), (5, -5)

hero (hero):

Then notice that 7 = 4 + 3 4 = 1 + 3 1 = -2 + 3 -2 = -5 + 3 And now notice this pattern \(r_n = r_{(n + 1)} + 3\)

hero (hero):

Can you see how to find the next term from here?